# Amorphous Solid Model of Vectorial Hopfield Neural Networks

### v1

@@ IV Discussion and Conclusions
In Fig. 5 we show the typical time evolution of the Lyapunov energy as a function of the number of updated steps, giving evidence of the monotonic convergence of the dynamics in the proposed neural network model.
05551010101515152020208-8- 86-6- 64-4- 42-2- 20Update stepEnergy
@@
Acknowledgments
AZ gratefully acknowledges funding from the US Army DEVCOM Army Research Office through contract nr. W911NF-22-2-0256 and from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”. In the making of this paper, we became aware of similar observations in a vectorial spin-glass model Nicoletti et al. (2025), that appeared shortly before this preprint.

### v2

@@ Acknowledgments
In Fig. 4, we show the eigenvalue distribution of the weight matrix $W_{ij}$ for the 3D networks proposed in this paper. The eigenvalue spectrum is obtained by direct numerical diagonalization of the weight matrix, Eq. (1), constructed for our networks. In Fig. 5 we show the typical time evolution of the Lyapunov energy as a function of the number of updated steps, giving evidence of the monotonic convergence of the dynamics in the proposed neural network model.
Figure 4: Evolution of the spectral density $\rho(\lambda)$ of the weight matrix $W_{ij}$ with different pattern capacities $\gamma=P/(Nd)$. Low density shows a concentrated band near zero with isolated positive eigenvalues, while high density exhibits broad continuous distributions, indicating pattern interference onset. The spectral density $\rho(\lambda)$ is obtained by direct numerical diagonalization of $W_{ij}$, as defined in Eq. (1), constructed for the 3D networks as in Fig. 1.
0551010151520208-86-64-42-20Update stepEnergy